noluvfoejaden:

I need help can y'all help me please or Evaluate the expression for c = –3. –15 − 3c

khloikamren0117:

c=−9/2c i think but I'm not for sure

luhivqqcherry:

-3(c+5)

noluvfoejaden:

@luhivqqcherry wrote:
-3(c+5)
@khloikamren0117 wrote:
c=−9/2c i think but I'm not for sure
it was -6

Onnagoroshiya:

-15+9 -6 right?

khloikamren0117:

@noluvfoejaden wrote:
@luhivqqcherry wrote:
-3(c+5)
@khloikamren0117 wrote:
c=−9/2c i think but I'm not for sure
it was -6
Oh, at least i tried lol

luhivqqcherry:

@khloikamren0117 wrote:
@noluvfoejaden wrote:
@luhivqqcherry wrote:
-3(c+5)
@khloikamren0117 wrote:
c=−9/2c i think but I'm not for sure
it was -6
Oh, at least i tried lol
samee.